OpenStudy (afia_arshad12):

Alex has twice as much money as Jennifer. Jennifer has $6 less than Shannon. Together they have $54. How much money does Alex have?

OpenStudy (zzr0ck3r):

when it says together, does it mean all three people?

OpenStudy (zzr0ck3r):

a=2j j=s-6 so s=j+6 a+j+s=54 2j + j + j+6 = 54 4j=48 j=12 a=24 s=18

OpenStudy (afia_arshad12):

Thanks @zzr0ck3r but it was a typo and It was supposed to be all of them. But thanks I got it now!

OpenStudy (zzr0ck3r):

I did all of them....